OpenStudy (nikki_skellington):

Polly uses a probability simulator to pull colored ribbons from a bag 60 times and to flip a coin 60 times. The results are shown in the tables below: Color of Ribbon Number of Times Pulled White 25 Green 14 Orange 21 Heads Tails 38 22 Using Polly's simulation, what is the probability of pulling an orange ribbon and the coin landing on tails? 308 over 3,600 330 over 3,600 462 over 3,600 36 over 60

jimthompson5910 (jim_thompson5910):

probability of orange = (# of orange)/(# of trials) = ???

OpenStudy (nikki_skellington):

there are 21 orange and 60 trials

OpenStudy (nikki_skellington):

so 0.35

jimthompson5910 (jim_thompson5910):

let's leave it as a fraction, so 21/60 which reduces to 7/20

OpenStudy (nikki_skellington):

ok

jimthompson5910 (jim_thompson5910):

probability of tails = (# of tails)/(# of coin flips) = ??

OpenStudy (nikki_skellington):

22 tails/60 trails right?

jimthompson5910 (jim_thompson5910):

22/60 which reduces to 11/30

OpenStudy (nikki_skellington):

ok..

jimthompson5910 (jim_thompson5910):

now multiply out 7/20 and 11/30

OpenStudy (nikki_skellington):

77/600

jimthompson5910 (jim_thompson5910):

hmm it looks your teacher didn't reduce the fractions (21/60)*(22/60) = (21*22)/(60*60) = 462/3600

jimthompson5910 (jim_thompson5910):

if you did reduce 462/3600, you would get 77/600

OpenStudy (nikki_skellington):

ok thx soo much!

jimthompson5910 (jim_thompson5910):

you're welcome